# Towards Physical Understanding in Video Generation: A 3D Point Regularization Approach

### v1

@@ Appendix A PointVid Dataset Generation
Starting with any video-text pair, we trim the videos to a fixed length $T$ and shape $H\times W$, extracting the first frame of the clipped video as the reference frame. We then apply semantic segmentation to the reference frame using Grounded-SAM2 [30, 29] to automatically generate masks for foreground objects. To streamline the process, we utilize a language model, LMDeploy [10], to infer the main moving objects in the scene based on the video’s caption. For example, the caption A man swinging a baseball bat in a studio results in {man, baseball bat}. This object-level label is subsequently fed to Grounded-SAM2 for prompt-based segmentation.
@@
Having obtained a reference frame and its segmentation, we apply SpaTracker [36] to track pixel movements accordingly. We use the same reference frame to query points during tracking so that the 3D positions align with the first frame’s pixel locations initially. We set downsample to 1 and gridsize to 80 in [36] to enable sparse tracking. For a 448$\times$256 image, we observe that sparse tracking at this resolution provides a sufficient number of pixels while remaining relatively efficient in terms of time. Increasing the resolution or using dense tracking could potentially enhance the results, but we leave this for future exploration. The model outputs point tracking data of shape $T\times N\times 3$, where $N$ is the number of sampled points.
Post-processing
To mitigate temporal noise caused by inaccurate tracking, we first apply a Kalman Filter to the 3D point cloud to estimate a global velocity and eliminate high-frequency noise. The result is then reshaped to $T\times H\times W\times 3$, where the spatial coordinates of pixels are treated as three channels, similar to RGB. To address the difference in shape between the $N$ tracked points and the video resolution, we apply the following rules: if a tracked pixel is not within the foreground mask, we discard its value (effectively setting all background pixels to zero). Otherwise, we assign its value to the corresponding pixel entry in the tensor. Since we use sparse tracking, only a subset of foreground pixels have values. We then apply KDTree searching on the sparse pixels based on their $uv$ coordinates in the first frame. For all untracked pixels within the foreground mask, we approximate their trajectory by interpolating their three nearest tracked pixels. The pseudocode for post-processing is summarized in 1.
@@
3:Initialize: $\mathcal{P}\leftarrow 0$triangleright Initialize all values to zero
@@
6:$\mathbf{P}_{1}\leftarrow\mathbf{P}_{0}\bigcap\mathbf{M}$triangleright Filter foreground pixels
7:$\mathbf{Tr}\leftarrow\text{KDTree}(\mathbf{P}_{1})$triangleright Use initial $uv$ for searching
@@
Appendix B Additional Qualitative Results
@@
We provide additional qualitative evaluations of our model compared to three other models: I2VGen-XL [41], SVD [6], DynamiCrafter [37]. These include both task-oriented videos (see Figure 7) and videos from non-specific categories (see Figure 8), such as humans, animals, and general objects. Our results demonstrate improved shape and motion consistency of objects, effectively minimizing common non-physical artifacts like object morphing.
Figure 7: Comparison on Task-oriented Videos.
Figure 8: Comparison on General Categories.
Figure 9: Ablation Studies on Different Components. We compare the results from (i) our full generation pipeline with (ii) model trained with RGB only, (iii) model trained with point augmentation only (no regularization), (iv) model without channel cross-attention, and (v) model trained without diffusion loss.
Appendix C Additional Ablation Study
We conduct ablation studies to evaluate the effectiveness of our training pipeline design. The results are visualized in Figure 9.
As discussed earlier, our two-stage design of point augmentation and regularization progressively enhances the video model. Compared to training the model with RGB videos only (Figure 9 (i)), our point augmentation alone (Figure 9 (ii)) injects 3D awareness into the model and improves its ability to perceive 3D shapes, as evidenced by more consistent object shapes in the videos. Our point regularization (Figure 9 (iii)) further improves quality by optimizing point generation and implicitly guiding RGB generation towards higher fidelity.
Although the video and point data are already pixel-aligned during generation, our channel-wise cross-attention mechanism is essential to ensure cross-dimensional and cross-modality alignment. Without the cross-attention (Figure 9 (iv)), 3D regularization on points is not correctly transferred to 2D pixels, leading to degraded video quality. Furthermore, in the regularization stage, we adopt a joint training strategy that combines regularization loss with diffusion loss. We observe that dropping diffusion loss (Figure 9 (v)) causes 3D information to dominate semantics, resulting in completely unusable output.
Appendix D User Study
@@
Apart from the quantitative evaluation using VBench [16] and VideoPhy [4], we additionally conduct a user study on our model in comparison to the base model I2VGen-XL [41] to assess human preference. We ask 10 labelers to compare the same testing batch used in quantitative evaluation (387 video clips), focusing on identifying non-physical artifacts and evaluating overall physical plausibility. The following 5 questions are designed for the evaluation:
@@
Given the evaluation questions, we ask users to compare our generated videos side by side with results from the base model and select the one that best fits each question. As shown in Table 2, our method demonstrates significant improvement in terms of physical plausibility by incorporating 3D awareness into the video.
Table 2: User Study Results. Our model demonstrates a significant improvement in physical plausibility, as assessed by human labelers. Here, Q2 asks the user to identify negative artifacts in the videos, while the other questions positively assess physical plausibility.

### v2

@@ A.1 PointVid Dataset Generation
Starting with any video-text pair, we trim the videos to a fixed length $T$ and shape $H\times W$, extracting the first frame of the clipped video as the reference frame. We then apply semantic segmentation to the reference frame using Grounded-SAM2 ren2024grounded ; ravi2024sam2segmentimages to automatically generate masks for foreground objects. To streamline the process, we utilize a language model, LMDeploy 2023lmdeploy , to infer the main moving objects in the scene based on the video’s caption. For example, the caption A man swinging a baseball bat in a studio results in {man, baseball bat}. This object-level label is subsequently fed to Grounded-SAM2 for prompt-based segmentation.
@@
Having obtained a reference frame and its segmentation, we apply SpaTracker xiao2024spatialtracker to track pixel movements accordingly. We use the same reference frame to query points during tracking so that the 3D positions align with the first frame’s pixel locations initially. We set downsample to 1 and gridsize to 80 in xiao2024spatialtracker to enable sparse tracking. We observe that sparse tracking at this resolution provides a sufficient number of pixels while remaining relatively efficient in terms of time. Increasing the resolution or using dense tracking could potentially enhance the results, but we leave this for future exploration. The model outputs point tracking data of shape $T\times N\times 3$, where $N$ is the number of sampled points.
@@
3:Initialize: $\mathcal{P}\leftarrow 0$ $\triangleright$ Initialize all values to zero
@@
[…]
@@
Post-processing
To mitigate temporal noise caused by inaccurate tracking, we first apply a Kalman Filter to the 3D point cloud to estimate a global velocity and eliminate high-frequency noise. The result is then reshaped to $T\times H\times W\times 3$, where the spatial coordinates of pixels are treated as three channels, similar to RGB. To address the difference in shape between the $N$ tracked points and the video resolution, we apply the following rules: if a tracked pixel is not within the foreground mask, we discard its value (effectively setting all background pixels to zero). Otherwise, we assign its value to the corresponding pixel entry in the tensor. Since we use sparse tracking, only a subset of foreground pixels have values. We then apply KDTree searching on the sparse pixels based on their $uv$ coordinates in the first frame. For all untracked pixels within the foreground mask, we approximate their trajectory by interpolating their three nearest tracked pixels. The pseudocode for post-processing is summarized in Algorithm 1.
A.2 Implementation Details
In this section, we provide more details on our experiments.
Dataset
We collect approximately 100K videos from chen2024panda and our internal dataset (each accounting for half) to form our PointVid dataset. The videos are processed to a resolution of 640 $\times$ 360 and cropped to a 2-second length (the exact number of frames may vary due to varying input frame rates, typically it is around 50 frames or more), and then paired with 3D point tracking using xiao2024spatialtracker .
Model Training
[…]
Joint Optimization
To effectively inject 3D-awareness into the video model without degrading the RGB modality (see subsection 4.3 for details), we adopt a joint optimization strategy combining diffusion and regularization losses. In our experiments, we observe that when the regularization loss significantly outweighs the diffusion loss (by orders of magnitude), it tends to suppress motion. Conversely, when the diffusion loss dominates training, non-physical artifacts reappear. In practice, we balance the two losses to maintain comparable magnitudes, which consistently reduces reconstruction misalignment and noticeably mitigates non-physical artifacts. To further preserve the motion magnitute, we apply the regularization loss less frequently (once every $k$ iterations, we use $k=5$). This adjustment preserves larger motion dynamics at the cost of requiring more training steps.
Computation Cost
For our UNet model trained on an A100 GPU with a batch size of 4, we report the additional computational cost as follows: incorporating 3D components and regularization losses increases training time ($\sim$9s per iteration, compared to $\sim$4.75s when fine-tuning on RGB video only) while maintaining comparable GPU memory usage (adding less than 2GB on top of $\sim$45GB).
A.3 Additional Qualitative Results
Additional Ablation Results
We provide additional qualitative results from our ablation studies in Figure 7. Our point augmentation and regularization approach progressively enhance the visual quality of the generated videos, while the channel cross-attention mechanism and diffusion loss are also critical for ensuring high-quality results.
Figure 7: Additional Ablation Results.
Additional Qualitative Comparisons
@@ A.3 Additional Qualitative Results
We provide additional qualitative evaluations of our model compared to other models: I2VGen-XL zhang2023i2vgen , SVD blattmann2023stablevideodiffusionscaling , DynamiCrafter xing2025dynamicrafter . These include both task-oriented videos (see Figure 8) and videos from non-specific categories (see Figure 9), such as humans, animals, and general objects. Our results demonstrate improved shape and motion consistency of objects, effectively minimizing common non-physical artifacts like object morphing.
Figure 8: Additioanl Comparison on Task-oriented Videos.
Figure 9: Additional Comparison on General Categories.
A.4 Additional Quantitative Results
While we use a fixed test set in all quantitative evaluations to ensure consistency, we also report the VBench huang2024vbench results of our UNet model using the official batch split provided by VBench. The results are summarized in Table 4. We observe similar outcomes as in the main comparison table (Table 1).
Table 4: Quantitative Results on VBench Test Batch. Our model outperforms the base model in most evaluation metrics.
Method SC $\uparrow$ BC $\uparrow$ MS $\uparrow$ AQ $\uparrow$ IQ $\uparrow$
Baseline 0.89703 0.94009 0.96993 0.52809 0.62961
Ours 0.96418 0.96722 0.98319 0.51754 0.65428
A.5 User Study
@@
Apart from the quantitative evaluation using VBench huang2024vbench and VideoPhy bansal2024videophy , we additionally conduct a user study on our UNet model in comparison to its base model I2VGen-XL zhang2023i2vgen to assess human preference on the models, focusing on identifying non-physical artifacts and evaluating overall visual plausibility. The following 5 questions are designed for evaluation:
@@
Given the evaluation questions, we ask 10 labelers to compare our generated videos side by side with results from the base model and select the one that best fits each question. We use the same testing batch used in quantitative evaluation. As shown in Table 5, our method demonstrates significant improvement in terms of physical plausibility by incorporating 3D awareness into the video.
Table 5: User Study Results. Our model demonstrates a significant improvement in physical plausibility, as assessed by human labelers. Here, Q2 asks the user to identify negative artifacts in the videos, while the other four questions positively assess physical plausibility.